Fermi National Laboratory
 

Operation Summary for April 12, 2010 - April 14, 2010

 

Monday April 12

The day shift began with the Tevatron in shot setup, the Recycler stash (416.5E10), and the Antiproton Source stack (20.6E10), with the Linac, Booster, and Main Injector operating normally, and with NuMI and MiniBooNE taking beam.

During the day shift, Operations established store 7737 at 9:20 AM with an average initial luminosity (AIL) of 377.4E30 (a new record). Operations established beam to MTest at 9:30 AM.

During the evening shift, Operations terminated store 7737 at 9:10 PM and put the TeV into shot setup. Operations established store 7738 at 10:22 PM with an AIL of 279E30. The F27 water skid tripped off causing minimal downtime to stacking. The water skid was reset.

Tuesday April 13

The midnight shift began with Operations monitoring store 7738 (187.09E30) and the stash (60.7E10), with Pbar stacking (20.1E10), and with NuMI, MiniBooNE and MTest taking beam.

During the midnight shift, Operations disabled beam to MTest at 4:01 AM. Operations terminated store 7738 at 5:21 AM. Operations prepared CDF, D0, Transfer Hall, F-Sector, A-E, Booster, Pbar, NuMI and MiniBooNE for scheduled access. The Main Injector and Recycler Ring (200.8E10) will not be accessed .

During the day shift, Feeder 46A was de-energized at 8:00 AM to find a ground fault and was re-energized at 3:00 PM. Experts began recovering the Tev cryogenic system once the feeder was energized. Recycler personnel conducted beam studies with the stash (200E10), when finished 100E10 remained.

During the evening shift, Operations established beam to the Linac dump at 5:04 PM, followed by beam to the Booster dump at 5:15 PM. Beam was established to MiniBooNE at 7:20 PM. Beam was established to Pbar at 9:14 PM, experts spent the rest of the shift investigating a low stack rate. Beam was established to NuMI at 9:45 PM. Tevatron cooldown continued through the entire shift and was complicated by a bad EVLH valve at F2.

Wednesday April 14

The midnight shift began with Operations monitoring the stash (110E10), with Pbar stacking (at a low rate), and with NuMI and MinBooNE taking beam.

During the midnight shift, experts continued to work on improving the PBar stacking rate. At 3:34 AM the Main Injector High Level RF tripped off on low temperature due to a loss of control to a 3 way valve that supplies cooling. Experts were called in, no beam to NuMI and Pbar. Tev F2 cooldown continued throughout the shift at a much slower than expected pace.

The Plans for Wednesday

The plan for today is to fix the MI high level RF problem, complete the Tev F2 cooldown and put in a Tev store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 04/14/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory